OpenStudy (anonymous):

solve the equation on the interval [0, 2pi) 8 sin^2x = -17 sinx - 9

OpenStudy (y2o2):

8sin²(x)+17sin(x)+9 = 0 [8sin(x) + 9][sin(x) + 1] = 0 sin(x) = -9/8 or sin(x) = -1

OpenStudy (anonymous):

At this point do I take the arcsin of -9/8 and of -1 to find every possible x in the domain?